X Company has the following costs for the year. Month January February March April May Units Produced Total Cost $ 710,000 665,000 670,000 680,000 695,000 10,600 6,300 6,800 7,300 9,100 June 9,600 698,000 July August September October November December 12,000 9,200 9,300 8,500 12,300 11,800 840,000 685,000 690,100 702,000 725,000 730,000 Using the High-Low method, compute the fixed cost per month.
Answer #1
Units Total cost
High level 12300 725000
Low level 6300 665000
Unit variable cost = (725000-665000)/(12300-6300)= $10
Fixed cost per month = 725000-(12300*10)= $602000